OpenStudy (anonymous):

I Have No Clue What This Is? The Area Of A Square Is Represented By (9z^2-12z+4)ft^2 . Which Expression Represents The Perimeter Of The Square?. A.(3z-2)ft B.(3z+2)ft C.(6z-4)ft D.(12z-8)ft ?

OpenStudy (anonymous):

it is a weird way to ask the question "what expression squared gives \(9z^2-12z+4\) ?"

OpenStudy (anonymous):

the are of a square is the square of the length of the side so for example if the are is \(25\) then the length of the side is \(5\) because \(5^2=25\) so really what they are after is do you recognize \(9z^2-12z+4\) as a perfect square ?

OpenStudy (anonymous):

*area

OpenStudy (anonymous):

@satellite73 Umm IDK? Is So Confusing!!

OpenStudy (anonymous):

So The Area Is The Whole Problem?.

OpenStudy (anonymous):

don't think very hard if you have \(9z^2-12z+4=(\text{something})^2\) it should be more or less clear what the "something" is

OpenStudy (anonymous):

the only way \(9x^2-12x+4\) will be a perfect square, is if it is the square of \(3z-2\)

OpenStudy (anonymous):

that is, check that \((3z-2)^2=9z^2-12z+4\)

OpenStudy (anonymous):

that means the length of the side is \(3z-2\) and so the perimeter is 4 times that, or \[4(3z-2)\]

OpenStudy (anonymous):

THANK Your A Real BIG Help! ((:

OpenStudy (anonymous):

yw